OpenStudy (anonymous):

Use the following problem to answer the question below: Given: 3(x – 4) = 33 Prove: x = 15

OpenStudy (anonymous):

Statments Reasons 3(x – 4) = 33 Given 3x – 12 = 33 A 3x – 12 + 12 = 33 + 12 B 3x + 0 = 45 C 3x = 45 D x = 15 Division Property of Equality What is the reason for A

OpenStudy (anonymous):

Distributive property of multiplication

OpenStudy (anonymous):

Can you also give me the reason for B and C @matricked Thanks

OpenStudy (anonymous):

Add 4 on each side. So, 3x now is equal to 37 or 3x = 37. Divide by 3 on each side. If we divide 3 on each side, we can see that you can't prove that x = 15 because it is not, x is actually equal to 12, not 15.